OpenStudy (anonymous):

solve for x: e^(x-7)/(e^-7)=28

OpenStudy (anonymous):

e^(x-7+7)=28 e^x=28 x=ln 28 x=3.33(2 decimal places)

OpenStudy (anonymous):

wait im confused about the signs of the 7s. wouldnt the denominator come up and equal to e^(x-7-7)?

OpenStudy (anonymous):

no because you had a \(-7\) in the denominator, so when it comes up it turns in to \(--7=7\)

OpenStudy (anonymous):

oh ok i wrote it wrong from the start ha. thank you very much